OpenStudy (anonymous):

Which fractions are equivalent to -4/9?

rebeccaxhawaii (rebeccaxhawaii):

multiply both the numerator and denominator by any number

rebeccaxhawaii (rebeccaxhawaii):

so if you multiply both by two that will give you an equivalent fraction

OpenStudy (anonymous):

Are the answers I picked correct?

OpenStudy (anonymous):

@rebeccaxhawaii ?

OpenStudy (anonymous):

@pooja195

OpenStudy (mathmale):

Look at one of your answers: \[\frac{ 16 }{ -36 }\]

OpenStudy (mathmale):

cAN YOU SHOW that this is the same fraction as -8/18? Can you show that -8/18 is the same fraction as -4/9 or 4/(-9)?

OpenStudy (anonymous):

ok, -8/18 divide both by 2= -4/9

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

What about 16/(-36)? what about -16/36? What is your conclusion about your own answers?

OpenStudy (anonymous):

But am I correct or not, that's all I want to know

OpenStudy (anonymous):

ugh

pooja195 (pooja195):

You are

OpenStudy (anonymous):

Ok thank you

pooja195 (pooja195):

yw